# Attrition-Aware Adaptation for Multi-Agent Patrolling

### v2

@@ Abstract
Multi-agent patrolling is a key problem in a variety of domains such as intrusion detection, area surveillance, and policing which involves repeated visits by a group of agents to specified points in an environment. While the problem is well-studied, most works do not provide performance guarantees and either do not consider agent attrition or impose significant communication requirements to enable adaptation. In this work, we present the Adaptive Heuristic-based Patrolling Algorithm, which is capable of adaptation to agent loss using minimal communication by taking advantage of Voronoi partitioning, and which meets guaranteed performance bounds. Additionally, we provide new centralized and distributed mathematical programming formulations of the patrolling problem, analyze the properties of Voronoi partitioning, and finally, show the value of our adaptive heuristic algorithm by comparison with various benchmark algorithms using a realistic simulation environment based on the Robot Operating System (ROS) 2.
@@ II Related Work
Recent work includes a method [20] for balancing agents’ priorities amongst important and unimportant observation points, generating routes for agents which take resource constraints into account. However, this method is not adaptive to disturbances such as agent attrition and unreliable communication. Another recent article by Kobayashi et al. presents a distributed patrol algorithm that ensures situation awareness of operators at a base station throughout the patrol [21], though it has the same drawbacks as [20].
@@ III-B Assumptions
  • An agent may only travel along the edges $E$ of the graph. If $\langle i,j\rangle\notin E\enspace\exists i,j\in V$, then an agent may not directly travel from node $i$ to node $j$.
@@ IV Mathematical Programming Approach
$\min_{x,y,u}$ $z=\frac{1}{n}\sum_{a=0}^{m}(u^{a}_{d_{a}}\sum_{i=0}^{n}y^{a}_{i})$ (IV)
s.t. $\sum_{\begin{subarray}{c}i=0\\ i\neq j\end{subarray}}^{n}x^{a}_{ij}=y^{a}_{j}\quad$ $\forall a\in A,\forall j\in\mathring{V}$ (1c)
$\sum_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}^{n}x^{a}_{ij}=y^{a}_{i}\quad$ $\forall a\in A,\forall i\in\mathring{V}$ (1f)
$u^{a}_{i}+c^{a}_{ij}-M(1-x^{a}_{ij})\leq u^{a}_{j}$ $\forall a\in A,\forall i\in V,$ (1g)
@@
$u^{a}_{o_{a}}=0\quad$ $\forall a\in A$ (1h)
$x^{a}_{d^{a}o^{a}}=1\quad$ $\forall a\in A$ (1i)
$\sum_{i\in D\setminus\{d^{a}\}}y^{a}_{i}=0\quad$ $\forall a\in A$ (1j)
$\sum_{a=0}^{m}y^{a}_{i}\geq 1\quad$ $\forall i\in V$ (1k)
@@
Constraint (1c) enforces that if agent $a$ visits node $j$, $a$ must have traversed some edge connected to $j$ (a “come-from” constraint). Constraint (1f) enforces that agent $a$ may only traverse a single edge immediately succeeding node $i$ (a “go-to” constraint). Constraint (1g) enforces that there are no subtours and records the visit time of agent $a$ to node $j$ as the visit time at node $i$ plus the travel cost (time) between nodes $i$ and $j$. This is similar to the “subtour elimination constraint” from the Miller-Tucker-Zemlin formulation of the TSP [26]. Constraint (1h) enforces that the origin node is visited at time 00. Constraint (1i) enforces that an agent $a$ must travel from its artificial destination node $d^{a}$ to its co-located origin node $o^{a}$, creating a cycle. Constraint (1j) enforces that an agent not visit the artificial destination nodes of other agents. Constraint (1k) enforces that all nodes must be visited.
@@
The objective function (IV) describes the average idleness time of all nodes. For each agent $a\in A$, we multiply the total time $u^{a}_{d^{a}}$ taken to complete a cycle by the number of nodes $\sum_{i\in V}y^{a}_{i}$ visited in that cycle. We define idleness as the time span between agent visits. Each of these nodes will then have the same idleness, since only one agent is assigned to each cycle. Summing the results for each agent $a$ provides the total idleness time of all nodes, which we divide by $n$ to find the average idleness time. Node idleness time is a commonly used metric in the literature for comparison of patrolling algorithms, as seen in e.g. [15][2][25]. By using this objective function, we can more easily compare our solution with existing approaches.
@@ IV-A Distributed Approach
$\min_{x^{a},u^{a}}$ $z^{a}=\frac{u^{a}_{d_{a}}}{n}\sum_{i=0}^{n}\hat{y}^{a}_{i}$ (IV-A)
s.t. $\sum_{\begin{subarray}{c}i=0\\ i\neq j\end{subarray}}^{n}x^{a}_{ij}=\hat{y}^{a}_{j}\quad$ $\forall j\in\mathring{V}$ (2c)
$\sum_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}^{n}x^{a}_{ij}=\hat{y}^{a}_{i}\quad$ $\forall i\in\mathring{V}$ (2f)
$u^{a}_{i}+c^{a}_{ij}-M(1-x^{a}_{ij})\leq u^{a}_{j}\quad$ $\forall i\in V,$ (2g)
@@ IV-B Agent Attrition
When an agent $a$ suffers attrition during execution, nodes assigned to that agent must be reassigned to other agents. In terms of the original MIP described in Section IV, this is equivalent to adding constraints $y^{a}_{i}=0\quad\forall i\in\mathring{V}$ and modifying (1i) to have a right-hand side of 00.
@@ V-A1 Agent Attrition & Adaptation
[…]
@@
where $c$ is not a neighboring agent of $\tilde{a}$. This implies that the path between $i$ and $c$ goes through a node $j$, with $a^{*}(j)=b$ and $b$ is a neighboring agent of $\tilde{a}$. For this node $j$, we also have
@@ VI Experimentation & Results
We experimentally compare our approach to several existing state-of-the-art dynamic patrolling methods using a realistic simulation environment that we developed for general multi-agent systems research.
VI-A Simulation Environment
@@
Separately, we have developed a multi-agent simulation and experimentation framework based on ROS 2, the Robot Operating System, which may be seen in Fig. 2. Our framework is capable of execution on real robots or may be used with a variety of compatible simulators. Using this multi-agent framework, we are able to write agent and experiment code which is common both to physical platforms and to a variety of simulators. ROS 2 provides us with features that are shared across these physical and simulated platforms including data visualization, communications introspection, runtime configuration, and a wide-ranging library of third-party modules for easy integration of new capabilities.
@@
Figure 2: Our MAS framework enables straightforward experimentation on multi-agent systems in both real and simulated environments.
We simulate sensor noise based on Gaussian distributions and require that each agent perform its own localization and navigation. This noise is the primary source of uncertainty in the simulation. Agents may collide or interference with each other, resulting in delay reaching goals.
@@ VI-B Benchmark Algorithms
We select four benchmark algorithms to compare against our own solution. Two are greedy algorithms that serve as a baseline: Greedy Bayesian Strategy (GBS) [15] and DTA-Greedy [2]. We also compare against the more sophisticated Concurrent Bayesian Learning Strategy (CBLS) [16] and DTA-Partitioning (DTAP or DTASSIPart) [2].
@@ VI-C Experimental Procedure
For each of the algorithms described in Section VI-B plus our own Algorithm 1 (AHPA), we perform a test using six agents in the 40-node “Cumberland” environment from [15]. A test lasts thirty minutes (1800 seconds). Each algorithm is tested twice: once with no agent attrition and once with attrition of single randomly-selected agents at the 300-second and 1300-second marks.
For each algorithm and each test, we execute three runs of 30 minutes each. Results are then averaged over the three runs to account for simulation noise or discrepancies. The test system is automated, and experimental monitoring and control is performed identically for each algorithm. All tests were performed on the same machine.
@@ VI-D Results
In analysis of the results, our primary focus is on the “average idleness” metric which we identified as our objective function in (IV). We look at this value over time to evaluate the performance of our algorithm.
Overall, our AHPA algorithm performed well in comparison with existing approaches and effectively addressed the problem statement in Section III. As seen in Fig. 3, the AHPA algorithm outperformed all others in non-attrition trials. We attribute this to AHPA’s partitioning of the environment and surmise via observation of experiments that many of the other algorithms suffered from either poor/overlapping partitioning (DTASSIPart) or from physical collisions amongst the agents (DTAGreedy, CBLS), which significantly impacts performance. AHPA’s solution resulted in non-overlapping partitions which improved allocations while also decreasing physical interference and collisions amongst the agents. As expected, use of AHPA resulted in a lower standard deviation of node idleness than other approaches, since each agent is assigned a tour which they then patrol continually. This may be seen in Fig. 3(b).
Figure 3: Performance of the algorithms over time, with no agent attrition. Note the excellent comparative performance of AHPA (in blue).
Figure 4: Performance of the algorithms over time, with two instances of agent attrition. Note the cumulative message difference between AHPA (blue) and the other algorithms.
In the attrition test, we removed a randomly-selected agent from the simulation at 300 seconds and another agent at 1300 seconds. The results of this are visible in Fig. 4(a). While AHPA did not recover faster than other algorithms, it required far fewer inter-agent messages than the other algorithms to do so. As seen in Fig. 4(c), it only required a single message in case of agent attrition. Contrast this with the many messages used by other algorithms, which can be problematic in communication-constrained environments.
We are especially surprised by the poor performance of the DTA-Partitioning algorithm (DTASSIPart in the graphs). Regardless of what we tried, we could not reproduce the high performance of that algorithm seen in [2]. We hypothesize that its poor performance was due to physical collisions and interference amongst the agents during mission execution. However, other algorithms did not suffer from this problem to the same extent.
@@
To demonstrate the AHPA performance bound described in Section V-B, we also compared AHPA with the optimal solution (computed using a commercial solver) for a toy problem with $n=6$ nodes and $m=2\dots 6$ agents. The results may be seen in Fig. 5, and demonstrate that useful performance bounds are provided.

### v3

@@ Abstract
Multi-agent patrolling is a key problem in a variety of domains such as intrusion detection, area surveillance, and policing which involves repeated visits by a group of agents to specified points in an environment. While the problem is well-studied, most works do not provide performance guarantees and either do not consider agent attrition or impose significant communication requirements to enable adaptation. In this work, we present the Adaptive Heuristic-based Patrolling Algorithm, which is capable of adaptation to agent loss using minimal communication by taking advantage of Voronoi partitioning, and which meets guaranteed performance bounds. Additionally, we provide new centralized and distributed mathematical programming formulations of the patrolling problem, analyze the properties of Voronoi partitioning, and finally, show the value of our adaptive heuristic algorithm by comparison with various benchmark algorithms using physical robots and simulation based on the Robot Operating System (ROS) 2.
@@ II Related Work
Recent work includes a method [20] for balancing agents’ priorities amongst important and unimportant observation points which generates routes for agents while taking resource constraints into account. However, this method is not adaptive to disturbances such as agent attrition and unreliable communication. Another recent article by Kobayashi et al. presents a distributed patrol algorithm that ensures situation awareness of operators at a base station throughout the patrol [21], though it has the same drawbacks as [20].
@@ III-B Assumptions
  • An agent may only travel along the edges $E$ of the graph. If $\langle i,j\rangle\notin E\enspace\exists i,j\in V$, then an agent may not directly travel from node $i$ to node $j$. An edge may be traversed by multiple agents simultaneously.
@@ IV Mathematical Programming Approach
$\min_{x,y,u}$ $z=\frac{1}{n}\sum_{a=0}^{m}(u^{a}_{d_{a}}\sum_{i=0}^{n}y^{a}_{i})$ (1)
s.t. $\sum_{\begin{subarray}{c}i=0\\ i\neq j\end{subarray}}^{n}x^{a}_{ij}=y^{a}_{j}\quad$ $\forall a\in A,\forall j\in\mathring{V}$ (1a)
$\sum_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}^{n}x^{a}_{ij}=y^{a}_{i}\quad$ $\forall a\in A,\forall i\in\mathring{V}$ (1b)
$u^{a}_{i}+c^{a}_{ij}-M(1-x^{a}_{ij})\leq u^{a}_{j}$ $\forall a\in A,\forall i\in V,$ (1c)
@@
$u^{a}_{o_{a}}=0\quad$ $\forall a\in A$ (1d)
$x^{a}_{d^{a}o^{a}}=1\quad$ $\forall a\in A$ (1e)
$\sum_{i\in D\setminus\{d^{a}\}}y^{a}_{i}=0\quad$ $\forall a\in A$ (1f)
$\sum_{a=0}^{m}y^{a}_{i}\geq 1\quad$ $\forall i\in V$ (1g)
@@
Constraint (1a) enforces that if agent $a$ visits node $j$, $a$ must have traversed some edge connected to $j$ (a “come-from” constraint). Constraint (1b) enforces that agent $a$ may only traverse a single edge immediately succeeding node $i$ (a “go-to” constraint). Constraint (1c) enforces that there are no subtours and records the visit time of agent $a$ to node $j$ as the visit time at node $i$ plus the travel cost (time) between nodes $i$ and $j$. This is similar to the “subtour elimination constraint” from the Miller-Tucker-Zemlin formulation of the TSP [26]. Constraint (1d) enforces that the origin node is visited at time 00. Constraint (1e) enforces that an agent $a$ must travel from its artificial destination node $d^{a}$ to its co-located origin node $o^{a}$, creating a cycle. Constraint (1f) enforces that an agent not visit the artificial destination nodes of other agents. Constraint (1g) enforces that all nodes must be visited.
@@
The objective function (1) describes the average idleness time of all nodes. For each agent $a\in A$, we multiply the total time $u^{a}_{d^{a}}$ taken to complete a cycle by the number of nodes $\sum_{i\in V}y^{a}_{i}$ visited in that cycle. We define idleness as the time span between agent visits. Each of these nodes will then have the same idleness, since only one agent is assigned to each cycle. Summing the results for each agent $a$ provides the total idleness time of all nodes, which we divide by $n$ to find the average idleness time. Node idleness time is a commonly used metric in the literature for comparison of patrolling algorithms, as seen in e.g. [15][2][25]. By using this objective function, we can more easily compare our solution with existing approaches.
@@ IV-A Distributed Approach
$\min_{x^{a},u^{a}}$ $z^{a}=\frac{u^{a}_{d_{a}}}{n}\sum_{i=0}^{n}\hat{y}^{a}_{i}$ (2)
s.t. $\sum_{\begin{subarray}{c}i=0\\ i\neq j\end{subarray}}^{n}x^{a}_{ij}=\hat{y}^{a}_{j}\quad$ $\forall j\in\mathring{V}$ (2a)
$\sum_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}^{n}x^{a}_{ij}=\hat{y}^{a}_{i}\quad$ $\forall i\in\mathring{V}$ (2b)
$u^{a}_{i}+c^{a}_{ij}-M(1-x^{a}_{ij})\leq u^{a}_{j}\quad$ $\forall i\in V,$ (2c)
@@ IV-B Agent Attrition
When an agent $a$ suffers attrition during execution, nodes assigned to that agent must be reassigned to other agents. In terms of the original MIP described in Section IV, this is equivalent to adding constraints $y^{a}_{i}=0\quad\forall i\in\mathring{V}$ and modifying (1e) to have a right-hand side of 00.
@@ V-A1 Agent Attrition & Adaptation
The agent $b\in A$ is considered a neighbor of agent $a\in A$ if for all nodes $i$ with $a^{*}(i)=a$ and $j$ with $a^{*}(j)=b$ and for all $t\in\mbox{path}(i,j)$, $a^{*}(t)=a$ or $a^{*}(t)=b$, where $\mbox{path}(i,j)$ refers to any shortest path between $i$ and $j$.
@@
where $c$ is not a neighboring agent of $\tilde{a}$. This implies that the path between $i$ and $c$ goes through a node $j$, with $a^{*}(j)=b$ and $b$ is a neighboring agent of $\tilde{a}$. For this node $j$, we also have that
@@ VI Experimentation & Results
We experimentally compare our approach to several existing state-of-the-art dynamic patrolling methods using both physical robots and a realistic simulation environment. Both physical and simulated experiments use the Grex framework, which we developed for general MAS research.
VI-A Experimental Environment
@@
Separately, we have developed a multi-agent simulation and experimentation framework based on ROS 2, the Robot Operating System, which may be seen in Fig. 2. Our framework, Grex, is capable of execution on real robots or may be used with a variety of compatible simulators. Using this multi-agent framework, we are able to write agent and experiment code which is common both to physical platforms and to a variety of simulators. ROS 2 provides us with features that are shared across these physical and simulated platforms including data visualization, communications introspection, runtime configuration, and a wide-ranging library of third-party modules for easy integration of new capabilities.
@@
Figure 2: We use the MAS framework, Grex, for straightforward experimentation in both simulated environments and with physical robots. At top-left are the TurtleBot3 robots which we use in physical experiments. At top-right, simulated robots patrol the “Cumberland” environment from [15].
We simulate sensor noise based on Gaussian distributions; this is the primary source of uncertainty in the simulation. Each agent performs its own localization and navigation. Agents may collide or interference with each other, resulting in delay reaching goals. For physical robots, real disturbances are at play, including sensor noise, localization uncertainty, communication losses, and robot collisions.
@@ VI-B Benchmark Algorithms
We select four benchmark algorithms to compare against our own solution. Two are greedy algorithms that serve as a baseline: Greedy Bayesian Strategy (GBS) [15] and DTA-Greedy [2]. We also compare against the more sophisticated Concurrent Bayesian Learning Strategy (CBLS) [16] and DTA-Partitioning (DTAP) [2].
@@ VI-C Experimental Procedure
For each of the algorithms described in Section VI-B plus our own Algorithm 1 (AHPA), we perform tests using six agents in both the 40-node “Cumberland” environment from [15] and our own 18-node “L440” environment, an empty conference room. A “Cumberland” test lasts thirty minutes (1800 seconds). Each algorithm is tested twice: once with no agent attrition and once with attrition of single randomly-selected agents at the 300-second and 1300-second marks. A test in the “L440” environment lasts for five minutes.
For each algorithm and each test, we execute three runs over which results are averaged to account for possible discrepancies. The test system is automated, and experimental monitoring and control is performed identically for each algorithm. All simulation tests were performed on the same machine.
@@ VI-D Results
In analysis of the results, our primary focus is on the “average idleness” metric which we identify as our objective function in (1). We look at this value over time to evaluate the performance of our algorithm.
Overall, our AHPA algorithm performs well in comparison with existing approaches and effectively addresses the problem statement in Section III. As seen in the left-hand column of Fig. 3, the AHPA algorithm outperforms all others in non-attrition trials. We attribute this to AHPA’s partitioning of the environment and surmise via observation of experiments that many of the other algorithms suffer from either poor/overlapping partitioning (DTAP) or from physical collisions amongst the agents (DTAG, CBLS), which significantly impacts performance. AHPA’s solution results in non-overlapping partitions which improves allocations while also decreasing physical interference and collisions amongst the agents. As expected, use of AHPA results in a lower standard deviation of node idleness than other approaches, since each agent is assigned a tour which they then patrol continually. This may be seen in the left-hand column of Fig. 3.
Figure 3: Performance of the algorithms over time in the simulated Cumberland environment originally used in [15]. At left, results with no agent attrition. Note the excellent comparative performance of AHPA (in blue). At right, with two instances of agent attrition. Note the cumulative message difference between AHPA and the other algorithms.
In the attrition test, we remove a randomly-selected agent from the simulation at 300 seconds and another agent at 1300 seconds. The results of this are visible in the right-hand column of Fig. 3. While AHPA does not recover faster than other algorithms, it requires far fewer inter-agent messages than the other algorithms to do so. As seen in Fig. 3, it only requires a single message in case of agent attrition. Contrast this with the many messages used by other algorithms, which can be problematic in communication-constrained environments.
Figure 4: Performance of AHPA and selected benchmark algorithms in real experiments. Note the stability of AHPA throughout. The DTAP and DTAG algorithms are not shown, as their performance was so poor that the graph became difficult to read.
Physical experiments demonstrate results consistent with those from simulation. AHPA exhibits stable operation throughout the test and outperforms the benchmark algorithms. However, we are surprised by the poor performance of the DTAP and DTAG algorithms. We cannot reproduce the high performance of those algorithms seen in [2]. In fact, DTAP and DTAG perform so poorly in our physical tests, with average idleness times nearing $252$ and $216$ seconds, respectively, that we exclude them from Fig. 4 for readability. We observe that their poor performance is due to physical collisions and interference amongst agents during mission execution. However, other algorithms do not suffer from this problem to the same extent.
@@
To demonstrate the AHPA performance bound described in Section V-B, we also compare AHPA with the optimal solution (computed using a commercial solver) for a toy problem with $n=6$ nodes and $m=2\dots 6$ agents. The results may be seen in Fig. 5, and demonstrate that useful performance bounds are provided.